OpenStudy (anonymous):

equation of the circle: (1,3), r=5 using the formula (x-h)^2+(y-k)^2

OpenStudy (anonymous):

(x-1)^2+(y-3)^2 = 25

OpenStudy (anonymous):

\[\Large h=1 \quad ,\quad k=3\] \[\Large (x-h)^2+(y-k)^2=r^2\] \[\Large [x-(1)]^2+[y-(3)]^2=5^2\] \[\Large (x-1)^2+(y-3)^2=25\]